CHARACTERIZATION OF STRUCTURAL MOTIFS FOR CO_{2} ACCOMMODATION BY IONIC SPECIES RELEVANT TO PHOTOELECTROCATALYSIS USING CRYOGENIC VIBRATIONAL PREDISSOCIATION SPECTROSCOPY

\underline{JOSEPH A. FOURNIER}, CHRISTOPHER J. JOHNSON, CONRAD T. WOLKE, AR-RON B. WOLK, CHRISTOPHER M. LEAVITT, KRISTEN J. BREEN, and MARK A. JOHNSON, DEPARTMENT OF CHEMISTRY, YALE UNIVERSITY, NEW HAVEN, CT 06520.

Two aspects of catalytic CO_{2} conversion to transportable fuels involve understanding the local interactions between CO_{2} and the solvent mediating the reduction and the coordination of CO_{2} onto the organometallic framework that executes the chemical transformation. To address why ionic liquids are emerging as the solvents of choice for the catalysis, we first explore how CO_{2} attaches to two archetypal ionic liquid components: imidazole (Im) and acetate (Ac^{−}). Im was seeded in a CO_{2} pulsed free jet expansion forming Im(CO_{2}^{−})(CO_{2})_{m} (m=1-2) clusters. Previous studies with pyridine revealed C-N bond formation between CO_{2}^{−} and pyridine with m=5 or 7, evidenced by the observation of a C-N stretch in the vibrational spectra. No such C-N bond formation is observed with Im; the CO_{2}^{−} appears to be hydrogen bonded to the Im N-H in an ion-molecule complex. Attachment to Ac^{−} was studied by predissociation of the Ac^{−}(CO_{2})_{m} (m=1-4) clusters and, for m=1-3, a single peak assigned to the CO_{2} asymmetric stretch is observed near that of bare CO_{2}, suggesting that the CO_{2} molecules are weakly bound adducts. However, the combination bands 2ν_{2} + ν_{3} and ν_{1} + ν_{3} in the 3600 − 3750cm^{−1} region reveal two distinct binding sites: a neutral site and a red-shifted, more perturbed site. For m=4, a red-shifted CO_{2} asymmetric stretch is observed, perhaps indicating the beginning of solvent mediated activation of a CO_{2} adduct.

Finally, the CO_{2} reduction catalyst Ni(cyclam)^{2+} has been investigated using an electrospray ionization source coupled to a cryogenically cooled ion trap. By adding CO_{2} to the He buffer gas used to collisionally cool the ions in the trap, we have been able to condense up to five CO_{2} molecules onto Ni(cyclam)^{2+} at a trap temperature of 100K. The cryogenic ion vibrational predissociation (CIVP) spectra show the CO_{2} molecules to be acting as neutral adducts. We are now working to develop a method to generate and isolate the reactive Ni(cyclam)^{+} and capture the bound-CO_{2} intermediate.